how many different 5 number combinations can there be using the numbers 1 threw 36
Answer by jim_thompson5910(35256)

There are 36*35*34*33*32 = 45239040 different permutations (where order matters)


So there are 45239040 /5! = 45239040/120 = 376992 different combinations (where order does NOT matter).


So the answer is 376992